OpenStudy (anonymous):

What is the binomial factor of y^15+8z^9?

OpenStudy (anonymous):

guess what? sum of two cubes. use \[a^3+b^3=(a+b)(a^2-ab+b^2)\] with \[a=y^5,b=2z^3\]

OpenStudy (anonymous):

No understanding.

OpenStudy (anonymous):

Just trying to help sister with study.

OpenStudy (anonymous):

you can read the formula i wrote yes? the left hand side is the sum of two cubes. the right hand side says how to factor it. \[y^{15}=(y^5)^3\] so put \[a=y^5\] \\[8z^0=(2z^3)^3\] so put \[b=(2z^3)\] and just place them in the formula on the right

OpenStudy (curry):

hey simba i need to talk to you

OpenStudy (curry):

simbasocks i need to ask you some questions about the power of one please